OpenStudy (anonymous):

MEDALS FOR BEST ANSWERS Solve for x: |x − 2| + 10 = 12 Solve for x: −3|2x + 6| = −12 Solve for x: |x + 2| + 16 = 14 Solve for x: −4|x + 5| = −16

OpenStudy (gorv):

|x-2|=2 x-2=+2 ............ x=4 and x-2=-2 .......... x=0

OpenStudy (gorv):

|2x+6|=4 2x+6=4 .........2x=-2 ......x=-1 2x+6=-4 ........... 2x=-10 .....x=-5

OpenStudy (anonymous):

modu;us function means u have to take positive and negative values.

OpenStudy (gorv):

|x+2|=-2 no solution mode always give positive values

OpenStudy (gorv):

|x+5|=4 x+5=4...........x=-1 x+5=-4........... x=-9

OpenStudy (gorv):

@lilman212